# The Impact of Auxiliary Patient Data on Automated Chest X-Ray Report Generation and How to Incorporate It

### v1

@@ 7 Limitations
Another limitation pertains to the completeness and quality of the patient data. Despite incorporating a wide range of data sources, the datasets still contain missing or incomplete information, which can affect model performance. For example, not all exams include a history section, and not all ED patient records have medicine administration details, leading to potential gaps in the data that the model can utilise. However, this reflects the nature of real patient records where issues of data quality and completeness are to be expected.
Our model’s architecture, while effective, has certain limitations. It struggles with large input sizes, especially when incorporating multiple prior exams, likely due to attention dilution. Future work should explore advanced attention mechanisms or hierarchical models to better manage large input sequences.
@@ 8 Ethical Considerations
Appendix A Prior exam embeddings
The images, findings section, and impression section from previous exams were considered. For prior exams, the time delta was positive, calculated by subtracting the time of the prior exam from the current exam. The images, findings section, and impression section from prior exams were given distinct source embeddings, separate from the current exam, to enhance differentiation. The comparison section from the current exam was also investigated, anticipating that references to prior exams in this section would prompt the decoder to reflect this in the generated report. We explored prior exams with a history size $h$ of up to three.
Appendix B Table column determination
@@
  • The column designations for the vital sign table were identical to the triage table, except for the rhythm column (e.g., ‘Normal Sinus Rhythm’), which was treated as a category column. The vital signs table also had no chief complaint column and the ‘charttime’ column was used as the event time.
  • For the medicine reconciliation table, the ‘intime’ from the ED stay table was used as the event time, as it pertains to the patient’s medication history prior to the ED stay. The name column was designated as a text column, while the gsn, ndc, etc_rn, and etccode columns were designated as category columns. The etcdescription column was not considered, as it is a descriprion of the etccode column.
  • For the medicine administration (pyxis) table, ‘charttime’ was used as the event time. The med_rn, name, gsn_rn, and gsn columns were all treated as category columns. The name column for the medicine reconciliation column did not have as high of a cardinality as the name column from the medicine reconciliation column, allowing it to be considered as a category column.
@@
Appendix C Experiment setup
C.1 Metrics
CheXbert-F1 (Smit et al., 2020), RadGraph-F1 Delbrouck et al. (2022), BLEU-4 (Papineni et al., 2001), and BERTScore-F1 (roberta-large_L17_no-idf_rescaled) (Zhang et al., 2020) have been found to correlate with radiologists’ assessment of reporting (Yu et al., 2023) and were a part of our evaluation. Additionally, we include CXR-BERT Boecking et al. (2022); Nicolson et al. (2024a), CIDEr (Vedantam et al., 2015), METEOR Banerjee and Lavie (2005), and ROUGE-L (Lin and Hovy, 2003) as part of our evaluation. CheXbert-F1, RadGraph-F1, and CXR-BERT were intended to capture the clinical semantic similarity between the generated and radiologist reports, while BERTscore-F1 was intended to capture general semantic similarity. Finally, CIDEr, METEOR ROUGE-L, and BLEU-4 were intended to capture the syntactic similarity between the generated and radiologist reports.
@@ C.2 Model
Our model is illustrated in Figure 2; following Nicolson et al. (2024b), we utilised UniFormer as the image encoder (in particular, the $384\times 384$ base model warm started with its token labelling fine-tuned checkpoint) Li et al. (2023). The image embeddings are formed by processing each image in the exam separately with the image encoder and then projecting its last hidden state to match the decoder’s hidden size using a learnable weight matrix. Each image was resized using bicubic interpolation so that its smallest side had a length of 384 and its largest side maintained the aspect ratio. Next, the resized image was cropped to a size of $\mathbb{R}^{3\times 384\times 384}$. The crop location was random during training and centred during testing. Following Elgendi et al. (2021), the image was rotated around its centre during training, where the angle of rotation was sampled from $\mathcal{U}{[{-5^{\circ},5^{\circ}}]}$. Finally, the image was standardised using the statistics provided with the UniFormer checkpoint. A maximum of five images per exam were used during training. If more were available, five were randomly sampled uniformly without replacement from the exam.
@@
Again following Nicolson et al. (2024b), we employed the Llama architecture for the decoder, which is notable for features such as its rotary positional encoding (RoPE), root mean square normalisation (RMSNorm), and SwiGLU activation function Touvron et al. (2023). A byte-level byte pair encoding tokenizer (Wang et al., 2020) was trained with a vocabulary size of 30 0003000030\,00030 000. It was trained on the findings, impression, indication, and history sections (not the comparison section) of the entire MIMIC-CXR training set, as well as the ‘pain’ and ‘chiefcomplaint’ columns from the triage table, the ‘name’ column of the medicine reconciliation table, and the ‘pain’ column from the vital signs table (from the entire MIMIC-IV-ED dataset). Newline, tab, repeated whitespaces, and leading and trailing whitespaces were removed from any text before tokenization.
@@
The hyperparameters of the Llama decoder were six hidden layers, a hidden size of 768, 12 attention heads per layer, and an intermediate size of 3 07230723\,0723 072. The maximum number of position embeddings was set to 2048 to accommodate all the patient data embeddings and the report tokens. The maximum number of tokens that could be generated was set to 256, which was also the limit for the radiologist reports during training. During testing, a beam size of four was utilised. The Llama decoder allows a custom attention mask to be provided in current implementations. [^4] https://huggingface.co/blog/poedator/4d-masks This enabled non-causal masking to be utilised for the prompt and causal masking for the report token embeddings, as shown in Figure 5. This ensured that the self-attention heads were able to attend to all of the patient data embeddings at each position.
@@
Figure 5: Attention mask for the decoder. Non-causal masking was used for the patient data embeddings and causal masking for the report token embeddings.
[…]
C.3 Training
@@
Three stages of training were performed. Each stage used AdamW (Loshchilov and Hutter, 2022) for mini-batch gradient descent optimisation, where training and evaluation was performed on a 94GB NVIDIA H100 GPU. The three stages were as follows:
@@
    Teacher forcing (TF) Williams and Zipser (1989) was performed on the MIMIC-CXR dataset with only the images for an exam as input, and exams that contained both a findings and impression section. This gave a training/validation split of 232 853232853232\,853232 853/1 83718371\,8371 837 images, 125 416125416125\,416125 416/991 exams, and 57 1015710157\,10157 101/436 patients. Training was performed with an initial learning rate of 5e-5, a mini-batch size of 8, a maximum of 32 epochs, and with float16 automatic mixed precision. All model parameters were trainable during this stage. The validation macro-averaged CheXbert-F1 was the monitored metric for checkpoint selection. This stage was necessary, as the language model struggled to generate reports from multiple sources without prior learning.
@@
    TF on the dataset described in Section 3 with the inputs described in Table 1. The training strategy was identical to the previous stage, except that a maximum of 16 epochs was performed, and the image encoder’s parameters were frozen (except for its projection). The models featured in Table 1 were trained using only the first two stages.
@@
    Reinforcement learning using self-critical sequence training (SCST) Rennie et al. (2017) with CXR-BERT and BERTScore as the reward (each weighted with 0.5) was performed in the final stage of training. The sample report for SCST was generated with top-k sampling ($k=50$). Training was performed with an initial learning rate of 5e-6, a mini-batch size of 32, a maximum of 24 epochs, and with float32 precision. The image encoder’s parameters were frozen during this stage (except for its projection). The validation BERTScore-F1 was the monitored metric for checkpoint selection, as it helped to select checkpoints less prone to repetitions. This stage of training was only applied to the best model from Table 1, ‘Images + effective sources (h = 0) - medicine reconciliation’, with the results presented in Table 2. This model had 161 185 728161185728161\,185\,728161 185 728 parameters.
C.4 Comparison Models
@@
Inputs RG CX CB BS
Images 26.24 28.36 57.17 24.33
Images + medicine administration 26.95 28.53 58.94 24.93
Appendix D Ancillary results
@@
In Figure 6, the F1-scores for each CheXbert label are shown. The ‘Images + effective sources ($h=0$) - medicine reconciliation’ model from Table 1 improves performance across all labels compared to the ‘Images’ model. This suggests that incorporating ancillary data from MIMIC-IV-ED and MIMIC-CXR provides a general improvement, rather than benefiting any specific pathology.
@@
Further improvements are seen when training the ‘Images + effective sources ($h=0$) - medicine reconciliation’ model with SCST (i.e., our model from Table 2) for most pathologies. However, there are performance decreases for ‘enlarged cardiomediastinum’, ‘lung lesion’, ‘pneumothorax’, and ‘fracture’. This might be due to these pathologies being underrepresented in the MIMIC-CXR dataset, leading the model to optimise for more common pathologies during SCST.

### v2

@@ 8 Limitations
This study currently lacks subjective evaluation by radiologists, which is essential for assessing the quality of generated reports. We plan to address this by evaluating with a private dataset and conducting radiologist-led assessments. To facilitate this, we are securing agreements and ethics approval for access to patient data and radiologist time. However, this process is extensive and beyond the scope of this study, and will instead be used to subjectively evaluate future models.
Another limitation pertains to the completeness and quality of the patient data. Despite incorporating a wide range of data sources, the datasets still contain missing or incomplete information, which can affect model performance. For example, not all exams include a history section, and not all ED patient records have administered medicines available, leading to potential gaps in the data that the model can utilise. However, this reflects the nature of real patient records where issues of data quality and completeness are to be expected.
Our model’s architecture, while effective, has certain limitations. It struggles with large input sizes, especially when incorporating multiple prior exams, likely due to attention dilution. It also at times struggles with supportive or confounding evidence from the auxiliary patient data, introducing false positive or false negative predictions. Future work should explore advanced attention mechanisms, hierarchical models, and LLMs to better manage large input sequences and to better balance auxiliary patient data evidence with imaging evidence.
@@ 9 Ethical Considerations
Appendix A Dataset Details
Each of the exams for the dataset described in Section 3 had one ED stay and triage row; 53% had at least one reconciled medicines row with up to 106 rows; 62% had at least one vital signs row with up to 69 rows; and 37% had at least one administered medicines row with up to 52 rows. Exams had an indication section 66% of the time with a maximum of 75 words, a history section 34% of the time with a maximum of 74 words, and a comparison section 97% of the time with a maximum of 129 words. Only one exam had both an indication and a history section.
Appendix B Prior Exam Embeddings
The images, findings section, and impression section from previous exams were considered. For prior exams, the time delta was positive, calculated by subtracting the time of the prior exam from the current exam. The images, findings section, and impression section from prior exams were given distinct source embeddings, separate from the current exam, to enhance differentiation. The comparison section from the current exam was also investigated, anticipating that references to prior exams in this section would prompt the decoder to reflect this in the generated report. We explored prior exams with a history size $h$ of up to three. Note that all exams from MIMIC-CXR were considered for the priors (train/validation/test 222 758222758222\,758222 758/1 80818081\,8081 808/3 26932693\,2693 269 exams), including those that did not occur during an ED stay and those that did not have a findings and/or impression section.
Appendix C Table Column Determination
@@
  • The column designations for the aperiodic vital signs table were identical to the triage table, except for the rhythm column (e.g., ‘Normal Sinus Rhythm’), which was treated as a category column. The aperiodic vital signs table also had no chief complaint column and the ‘charttime’ column was used as the event time.
  • For the reconciled medicines table, the ‘intime’ from the ED stay table was used as the event time, as it pertains to the patient’s medicine history prior to the ED stay. The name column was designated as a text column, while the gsn, ndc, etc_rn, and etccode columns were designated as category columns. The etcdescription column was not considered, as it is a descriprion of the etccode column.
  • For the administered medicines (pyxis) table, ‘charttime’ was used as the event time. The med_rn, name, gsn_rn, and gsn columns were all treated as category columns. The name column for the administered medicines column did not have as high of a cardinality as the name column from the reconciled medicines column, allowing it to be considered as a category column.
@@
Appendix D Experiment Setup
D.1 Metrics
GREEN Ostmeier et al. (2024), CheXbert-F1 (Smit et al., 2020), RadGraph-F1 Delbrouck et al. (2022), BLEU-4 (Papineni et al., 2001), and BERTScore-F1 (roberta-large_L17_no-idf_rescaled) (Zhang et al., 2020) have been found to correlate with radiologists’ assessment of reporting (Yu et al., 2023; Ostmeier et al., 2024) and were a part of our evaluation. Additionally, we include CXR-BERT Boecking et al. (2022); Nicolson et al. (2024a), and ROUGE-L (Lin and Hovy, 2003). GREEN, CheXbert-F1, RadGraph-F1, and CXR-BERT were intended to capture the clinical semantic similarity between the generated and radiologist reports, while BERTscore-F1 was intended to capture general semantic similarity. Finally, ROUGE-L and BLEU-4 were intended to capture the syntactic similarity between the generated and radiologist reports. We also propose a new metric that measures n-gram repetition rate, namely the absence of repeated n-grams (ARN). It is calculated as:
$\text{ARN}=\begin{cases}1.0&\text{if }L<n,\\ 1.0-\frac{\sum_{i=1}^{M}(\text{Count}(g_{i})-1)}{M}&\text{if }L\geq n,\end{cases}$ (1)
where $L$ is the total number of tokens in the generated report, $n$ is the $n$-gram size, $M=L-n+1$ is the total number of $n$-grams in the report, $g_{i}$ is the $i^{th}$ unique $n$-gram in the report, $\textrm{Count}(g_{i})$ is the $n$-gram frequency in the report. The tokenizer described in Appendix D.2 was used with an $n$-gram size of three.
@@ D.2 Model
Our model is illustrated in Figure 2; following Nicolson et al. (2024b), we utilised UniFormer as the image encoder (in particular, the $384\times 384$ base model warm started with its token labelling fine-tuned checkpoint) Li et al. (2023). The image embeddings are formed by processing each image in the exam separately with the image encoder and then projecting its last hidden state to match the decoder’s hidden size using a learnable weight matrix. Each image was resized using bicubic interpolation so that its smallest side had a length of 384 and its largest side maintained the aspect ratio. Next, the resized image was cropped to a size of $\mathbb{R}^{3\times 384\times 384}$. The crop location was random during training and centred during testing. Following Elgendi et al. (2021), the image was rotated around its centre during training, where the angle of rotation was sampled from $\mathcal{U}{[{-5^{\circ},5^{\circ}}]}$. Finally, the image was standardised using the statistics provided with the UniFormer checkpoint. A maximum of five images per exam were used during training. If more were available, five were randomly sampled uniformly without replacement from the exam for each epoch.
@@
Again following Nicolson et al. (2024b), we employed the Llama architecture for the decoder, which is notable for features such as its rotary positional encoding (RoPE), root mean square normalisation (RMSNorm), and SwiGLU activation function Touvron et al. (2023). A byte-level byte pair encoding tokenizer (Wang et al., 2020) was trained with a vocabulary size of 30 0003000030\,00030 000. It was trained on the findings, impression, indication, and history sections (not the comparison section) of the entire MIMIC-CXR training set, as well as the ‘pain’ and ‘chiefcomplaint’ columns from the triage table, the ‘name’ column of the reconciled medicines table, and the ‘pain’ column from the vital signs table (from the entire MIMIC-IV-ED dataset). Newline, tab, repeated whitespaces, and leading and trailing whitespaces were removed from any text before tokenization.
@@
The hyperparameters of the Llama decoder were six hidden layers, a hidden size of 768, 12 attention heads per layer, and an intermediate size of 3 07230723\,0723 072. The maximum number of position embeddings was set to 2 04820482\,0482 048 to accommodate all the patient data embeddings and the report tokens. The maximum number of tokens that could be generated was set to 256, which was also the limit for the radiologist reports during training. During testing, a beam size of four was utilised. The Llama decoder allows a custom attention mask to be provided in current implementations. [^3] https://huggingface.co/blog/poedator/4d-masks This enabled non-causal masking to be utilised for the prompt and causal masking for the report token embeddings, as shown in Figure D.1. This ensured that the self-attention heads were able to attend to all of the patient data embeddings at each position.
@@
Figure D.1: Attention mask for the decoder. Non-causal masking was used for the patient data embeddings and causal masking for the report token embeddings.
D.3 Training
@@
Three stages of training were performed. Each stage used AdamW (Loshchilov and Hutter, 2022) for mini-batch gradient descent optimisation and gradient clipping with a maximum norm of 1.0 to prevent exploding gradients and maintain training stability. Training and evaluation was performed on a 94GB NVIDIA H100 GPU. The three stages were as follows:
@@
    Teacher forcing (TF) Williams and Zipser (1989) was performed on the MIMIC-CXR dataset with only the images for each exam as input, and exams that contained both a findings and impression section. This gave a training/validation split of 232 855232855232\,855232 855/1 83718371\,8371 837 images, 125 417125417125\,417125 417/991 exams, and 57 1025710257\,10257 102/436 patients. Training was performed with an initial learning rate of 5e-5, a mini-batch size of 8, a maximum of 32 epochs, and with float16 automatic mixed precision. All model parameters were trainable during this stage. The validation macro-averaged CheXbert-F1 was the monitored metric for checkpoint selection. This stage was necessary, as the language model struggled to generate reports from multiple patient data sources without prior learning.
@@
    TF was used in the second stage of training, with the MIMIC-CXR & MIMIC-IV-ED dataset described in Section 3 with the inputs described in Table 1. The training strategy was identical to the previous stage, except that a maximum of 16 epochs was performed, and the image encoder’s parameters were frozen (except for its projection). The models featured in Table 1 were trained using only the first two stages.
@@
    Reinforcement learning using self-critical sequence training (SCST) Rennie et al. (2017) was performed with the rewards described in Appendix E in the final stage of training. The sample report for SCST was generated with top-k sampling ($k=50$). Training was performed with an initial learning rate of 5e-6, a mini-batch size of 32, a maximum of 32 epochs, and with float32 precision. A warmup phase of 5 00050005\,0005 000 training steps was used for the learning rate, linearly increasing from zero. The image encoder’s parameters were frozen during this stage (except for its projection). The validation BERTScore-F1 was the monitored metric for checkpoint selection. This stage of training was only applied to the best model from Table 1, ‘Images + effective sources (h = 0)’, with the results presented in Table 2.
D.4 Comparison Models
@@
Appendix E Reinforcement Learning Rewards
The separate reward per section was calculated as:
$\begin{split}r_{s}(\hat{\textbf{w}}_{f},\textbf{w}_{f},\hat{\textbf{w}}_{i},\textbf{w}_{i})=&\alpha_{1}\cdot r_{f}(\hat{\textbf{w}}_{f},\textbf{w}_{f})+\\ &\alpha_{2}\cdot r_{i}(\hat{\textbf{w}}_{i},\textbf{w}_{i}),\\ \end{split}$ (2)
where $r_{s}(\cdot)$ is the composite reward for the sections of the report, $r_{f}(\cdot)$ is the reward for the findings section, and $r_{i}(\cdot)$ is the reward for the impression section, $\hat{\textbf{w}}_{f}$ is the generated findings section, $\textbf{w}_{f}$ is the radiologist findings section, $\hat{\textbf{w}}_{i}$ is the generated impression section, $\textbf{w}_{i}$ is the radiologist impression section, and $\alpha_{1}$ and $\alpha_{2}$ are weights. Normally, $r_{r}(\hat{\textbf{w}}_{r},\textbf{w}_{r})$ is calculated, where $\hat{\textbf{w}}_{r}$ and $\textbf{w}_{r}$ are the generated and radiologist reports, which include both the findings and impression sections.
The reward $r_{f}(\cdot)$, $r_{i}(\cdot)$, or $r_{r}(\cdot)$ is calculated as:
$\begin{split}r(\hat{\textbf{w}},\textbf{w})=&\lambda_{1}\cdot\textrm{CXR-BERT}(\hat{\textbf{w}},\textbf{w})+\\ &\lambda_{2}\cdot\textrm{BERTScore}(\hat{\textbf{w}},\textbf{w})+\\ &\lambda_{3}\cdot\textrm{ARN}(\hat{\textbf{w}},\textbf{w}),\\ \end{split}$ (3)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights. For ‘Images + effective source ($h=0$) + RL with CXR-BERT + BERTScore reward’, $\lambda_{1}=0.5$, $\lambda_{2}=0.5$, and $\lambda_{3}=0.0$. For ‘Images + effective source ($h=0$) + RL with CXR-BERT + BERTScore reward per section’, $\alpha_{1}=0.75$, $\alpha_{2}=0.25$, $\lambda_{1}=0.5$, $\lambda_{2}=0.5$, and $\lambda_{3}=0.0$. A higher weight was used for the findings section, as it is longer on average than the impression section. For ‘Images + effective source ($h=0$) + RL with CXR-BERT + BERTScore + ARN reward per section’, $\alpha_{1}=0.75$, $\alpha_{2}=0.25$, $\lambda_{1}=0.45$, $\lambda_{2}=0.45$, and $\lambda_{3}=0.1$. Only a weak contribution of the ARN was required to prevent repetitions.
The improvement that separating the reward per section has on the findings section is negligible, as seen in Table 2. However, separating the reward per section improves the scores for the impression section, as shown in Table E.1. Separating the reward likely enables the model to better optimise for the concise and summarised nature of the impression section, which was previously overshadowed by the dominance of the findings section’s requirement for comprehensive detail when both were jointly considered.
@@ Appendix E Reinforcement Learning Rewards
Model RG CX CB G BS R-L B4 ARN
Images + effective sources ($h=0$) 20.21 26.81 57.61 28.71 27.90 25.02 4.77 99.59
   + RL (CXR-BERT + BERTScore reward) 23.96 28.07 62.85 30.58 31.58 28.48 7.84 99.89
    + reward per section 24.89 31.08 71.12 30.89 36.27 30.27 6.70 99.33
     + ARN reward 24.87 32.88 71.12 32.14 36.31 30.61 6.84 99.83
Figure E.1: The mean pairwise difference GREEN score for the generated report (findings and impression sections) of each exam from the test set between 10 training runs of the “Images” model and the “Images + effective sources (h=0)” model. This illustrates the performance change (increase or decrease) over the exams resulting from incorporating auxiliary patient data for different CheXpert labels. $\Delta$, $m$ and $n$ are the number of training runs for each model ($m=n=10$) and $s$ is the GREEN score for one of the models. The subplots are sorted in descending order based on the ratio of $\textrm{AUC}(\Delta>0)$ to $\textrm{AUC}(\Delta<0)$.
[…]
Appendix F Ancillary Results
@@
In Figure E.2, the F1-scores for each CheXbert label are shown. The ‘Images + effective sources ($h=0$)’ model from Table 1 attained a higher score than the ‘Images’ model for 11 of the 14 labels. This suggests that incorporating auxiliary patient data from MIMIC-IV-ED and MIMIC-CXR provides a general improvement, rather than benefiting any specific pathology.
@@
Further improvements can be seen for most labels when reinforcement learning (RL) is used (i.e., our model from Table 2). However, there are performance decreases for ‘enlarged cardiomediastinum’, ‘pneumothorax’, and ‘fracture’. This might be due to these pathologies being underrepresented in the MIMIC-CXR dataset, leading the model to optimise for more common pathologies during reinforcement learning.